In the last year, Bakers Score's Donuts and Sweets has sold, on average, 1,000 ''Fudgie'' brownies per week. This week they introduced fat-free brownies and sold 200 at a price of $3.84 each. They noticed the cannibalization rate of fat-free brownies on Fudgie brownies was 46%. Bakers Score's sells their "Fudgie" brownies for $3.10, with a variable cost of $1.40. The new Fat-Free Brownies have a variable cost of $1.64. What is the effective unit margin ($) including the impact of cannibalization for Fat-Free Brownies, assuming the original estimate of cannibalization.
Answer #1

Before the introduction of the fat free brownie,

sales per week = 1000

Contribution margin = sale price - variable cost = 3.1-1.4=1.7

Total margin = 1.7*1000=1700$

With introduction of fat free brownie,

sale per week of fat free brownie = 200

Contribution margin = sale price - variable cost = 3.84-1.64=2.2

Total margin from fat free brownie = 2.2*200=440$

sale per week of normal brownie = (1- cannibalisation rate)* normal sales per week = (1-46%)*1000=540

Contribution margin = sale price - variable cost = 3.1-1.4=1.7

Total margin from normal brownie = 1.7*540=918$

Total margin after introduction of fat free brownie = 440$+918$=$1358

Effective unit margin = total margin / no of units = 1358/ (200+540)=$1.835